# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ Abstract
We propose new vertex operators, both the type I and the type II dual, of the elliptic quantum toroidal algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ by combining representations of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ and the notions of the elliptic stable envelopes for the instanton moduli space ${\cal M}(n,r)$. The vertex operators reproduce the shuffle product formula of the elliptic stable envelopes by their composition. We also show that the vacuum expectation value of a composition of the vertex operators gives a correct formula of the $\mathrm{K}$-theoretic vertex function for ${\cal M}(n,r)$. We then derive exchange relations among the vertex operators and construct a $L$-operator satisfying the $RLL=LLR^{*}$ relation with $R$ and $R^{*}$ being elliptic dynamical instanton $R$-matrices defined as transition matrices of the elliptic stable envelopes. Assuming a universal form of $L$, we define a comultiplication $\Delta$ in terms of it. It turns out that the new vertex operators are intertwining operators of the $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$-modules w.r.t $\Delta$.
1 Introduction
@@
The aim of this paper is to present a new formulation of the vertex operators of the elliptic quantum toroidal algebra (EQTA) $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ by combining its representations and the notions of the elliptic stable envelopes for the instanton moduli space ${\cal M}(n,r)$.
The EQTA $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ is an elliptic quantum group associated with the toroidal algebra of type ${\mathfrak{gl}}_{1}$ formulated in [32]. Here parameters $t_{1}$ and $t_{2}$ correspond to $q^{-1}$ and $t$ in [32], respectively. It is defined by generators and relations in the same scheme as the elliptic quantum group (EQG) $U_{q,p}(\widehat{\mathfrak{g}})$ associated with the affine Lie algebra $\widehat{\mathfrak{g}}$[23, 19, 7, 31] and equipped with a Hopf algebroid structure given by the Drinfeld comultiplication, i.e. a comultiplication formula in terms of the Drinfeld generators or their generating functions called the elliptic currents. Two types of vertex operators, type I and type II dual, have been constructed as intertwining operators of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$-modules and shown to give a realization of the generating function of the affine quiver $W$-algebra associated with the Jordan quiver $\Gamma(\widehat{A}_{0})$[20]. At the same time, the elliptic currents of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ gives the screening currents of it. In addition, it has been shown that these vertex operators realize the refined topological vertices[15, 16] relevant to calculate the instanton partition functions of the 5d and 6d lift of the 4d ${\cal N}=2^{*}$ gauge theory[39, 40]. However their relations to the elliptic stable envelopes and to the vertex functions of the corresponding quiver variety were missing.
@@
The notion of stable envelopes was initiated in [35, 41]. They provide a good basis of equivariant cohomology and $\mathrm{K}$-theory of Nakajima quiver varieties $X$ [37, 38] and has nice applications to enumerative geometry, geometric representation theory and quantum integrable systems. In particular, the transition matrices between stable envelopes defined for different chambers give the $R$-matrices satisfying the Yang-Baxter equation (YBE). Regarding such $R$-matrix as universal one i.e. a $L$-operator which generate a quantum group, the transition property can be understood as a Gauss decomposition of the $L$-operator in terms of the stable envelopes. Hence the stable envelopes provide a new geometric formulation of quantum groups as well as quantum integrable systems associated with the quiver varieties. The elliptic version of stable envelopes (ESE) [1] for the equivariant elliptic cohomology $\mathrm{E}_{T}(X)$ is more fascinating. They depend on both the equivariant and the Kähler parameters and provide the elliptic dynamical $R$-matrix as their transition matrix. There these two parameters play a role of the spectral and the dynamical parameters, respectively.
@@
An typical example is the case $X=T^{*}fl$, the cotangent bundle to the partial flag variety. The ESE’s for $\mathrm{E}_{T}(T^{*}fl)$ can be identified with the elliptic weight functions of the $A$ type appearing in the elliptic hypergeometric integral solutions of the elliptic $q$-KZ equation[28, 29, 43] and their transition matrices coincide with the known elliptic dynamical $R$-matrices of type $A$[17, 12, 18, 24]. In [28, 29], such identification was carried out by using the vertex operators defined as intertwining operators of the $U_{q,p}(\widehat{\mathfrak{sl}}_{N})$-modules w.r.t the standard comultiplication given in terms of the $L$-operator instead of the Drinfeld comultiplication. Basically such vertex operators are operator valued integrals, the so called screened vertex operators, possessing a basic elliptic weight function i.e. elliptic stable envelope, as an integration kernel. Then a composition of them produces more general ESE corresponding to a general configulation in $T^{*}fl$ via the shuffle product formula [29]. Such structure can indeed be regarded as an inheritance from the $L$-operator of $U_{q,p}(\widehat{\mathfrak{sl}}_{N})$, whose Gauss components are given by the so called half currents, i.e. integrals of the elliptic currents of $U_{q,p}(\widehat{\mathfrak{sl}}_{N})$ with the same basic elliptic weight function as an integration kernel[22]. This is consistent to a picture that the stable envelope gives a half of the Gauss decomposition of the $L$-operator.
@@
In addition, for the same case $X=T^{*}fl$, it was shown that the vertex operators yield a ${\mathrm{K}}$-theoretic vertex functions[41, 42, 33] as a vacuum expectation value of a composition of the vertex operators[30]. The vertex function is a generating function of the counting of quasi maps from $X$ to ${\mathbb{P}}^{1}$. It is given by a certain $q$-hypergeometric integral containing the elliptic weight function inside the integral and satisfies the $q$-KZ equation. The monodromy matrix of it is then given by a transition matrix of the weight functions. In [1], such elliptic weight functions are called the pole subtraction elliptic stable envelopes.
Motivated by such nice relationship between representation theory of $U_{q,p}(\widehat{\mathfrak{g}})$ and the elliptic stable envelopes, we in this paper propose new vertex operators of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ by combining representations of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$[32] and the elliptic stable envelopes for the equivariant elliptic cohomology $\mathrm{E}_{T}({\cal M}(n,r))$ of the instanton moduli space ${\cal M}(n,r)$ constructed in [44, 4]. We construct both the type I and the type II dual vertex operators as screened vertex operators whose integration kernel is given by the ESE for $\mathrm{E}_{T}({\cal M}(n,r))$. We then make several checks on their consistency. One of them is the shuffle product formula of ESE’s. We derive it by considering a composition of the vertex operators and check its consistency with the one obtained geometrically in [2]. We also construct $\mathrm{K}$-theoretic vertex functions for ${\cal M}(n,r)$ as expectation value of compositions of the vertex operators and show their coincidence with the formula obtained in [45, 4]. Moreover, thus obtained vertex operators, in particular for the Hilbert scheme $\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$, allow us to construct a $L$-operator $L^{+}(u)$ on a certain representation, which satisfies the $RLL=LLR^{*}$ relation. Here $R$ and $R^{*}$ are elliptic dynamical $R$-matrices given as transition matrices of ESE’s for $\mathrm{E}_{T}(\mathrm{Hilb}^{n}({\mathbb{C}}^{2}))$ appearing in the type I and the type II dual vertex operators, respectively. Hence they should be the elliptic instanton $R$-matrices. We then consider a universal form ${\cal L}^{+}(u)$ of $L^{+}(u)$ and show that these vertex operators are intertwining operators of the $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$-modules w.r.t the standard comultiplication defined by ${\cal L}^{+}(u)$.
@@
This paper is organized as follows. Sec.2 is a review of the elliptic quantum toroidal algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ including the Hopf algebroid structure and representations of the level $(1,N)$ and $(0,1)$. In Sec.3, we summarize known results on the ESE for $\mathrm{E}_{T}({\cal M}(n,r))$. In Sec.4, we propose a new construction of the type I and the type II dual vertex operators of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ and check the shuffle product formula of ESE’s. Sec.5 is devoted to a study of the $\mathrm{K}$-theoretic vertex functions for ${\cal M}(n,r)$. In Sec.6, we derive exchange relations of the vertex operators. Finally, in Sec.7, we construct a $L$-operator $L^{+}(u)$, which satisfies the $RLL=LLR^{*}$ relation. The exchange relations between the vertex operators and $L^{+}(u)$ are given. Defining a new comultiplication $\Delta$, we show that our vertex operators are intertwining operators of the $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$-modules w.r.t. $\Delta$. For comparison, we list in Appendix A a summary of the vertex operators w.r.t. the Drinfeld comultiplication $\Delta^{D}$ obtained in [32].
2 Elliptic Quantum Toroidal Algebra ${U}_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$
@@
We review the elliptic quantum toriodal algebra $\mathcal{U}_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ introduced in [32]. Parameters $t_{1},t_{2},\hbar=t_{1}t_{2}$ in this paper correspond to $q^{-1},t,t/q$ in [32], respectively. We assume $p,t_{1},t_{2}$ being generic complex number and $|p|<1$.
2.1 Definition
@@
Let us consider the Heisenberg algebras generated by $c,\ \Lambda_{0},\ c^{\perp},\ \Lambda_{0}^{\perp},\ P,\ Q$ satisfying the commutation relations
$[c,\Lambda_{0}]=1=[c^{\perp},\Lambda_{0}^{\perp}]=[P,Q],$ (2.1)
@@
the others are zero. We set $\gamma=\hbar^{c/2},\ C=\hbar^{c^{\perp}/2},\ \mathfrak{z}^{*}=\hbar^{P}$ and $\mathfrak{z}=\mathfrak{z}^{*}C^{2}=\hbar^{P+c^{\perp}}$. We call $\mathfrak{z}^{*}$ the dynamical parameter. Let ${\mathbb{F}}$ be the field of meromorphic functions of $\mathfrak{z}$ and $\mathfrak{z}^{*}$. We have
$g(\mathfrak{z},\mathfrak{z}^{*})e^{\Lambda_{0}^{\perp}-Q}=e^{\Lambda_{0}^{\perp}-Q}g(\mathfrak{z},\mathfrak{z}^{*}\hbar^{-1}),\quad g(\mathfrak{z},\mathfrak{z}^{*})e^{-\Lambda_{0}^{\perp}}=e^{-\Lambda_{0}^{\perp}}g(\mathfrak{z}\hbar^{-1},\mathfrak{z}^{*})\qquad\forall g(\mathfrak{z},\mathfrak{z}^{*})\in{\mathbb{F}}.$
Set
@@
The elliptic quantum toroidal algebra $\mathcal{U}=U_{t_{1},t_{2},p}({{\mathfrak{gl}}}_{1,tor})$ is a topological associative algebra over ${\mathbb{F}}[[p]]$ generated by $\alpha_{m},x^{\pm}_{n}$, $(m\in{\mathbb{Z}}\backslash\{0\},n\in{\mathbb{Z}})$ and $C,\gamma^{1/2}$. Let $x^{\pm}(z),\psi^{\pm}(z)$ be the following generating functions [^1] Our $C$ and $x^{\pm}(z)$ here are $\psi_{0}^{-}=(\psi_{0}^{+})^{-1}$ and $x^{\pm}(\gamma^{1/2}z)$ in [32], respectively. .
@@
$\psi^{+}(z):=C^{-1}\exp\left(-\sum_{m>0}\frac{p^{m}}{1-p^{m}}\alpha_{-m}(\gamma^{-1/2}z)^{m}\right)\exp\left(\sum_{m>0}\frac{1}{1-p^{m}}\alpha_{m}(\gamma^{-1/2}z)^{-m}\right),$
$\psi^{-}(z):=C\exp\left(-\sum_{m>0}\frac{1}{1-p^{m}}\alpha_{-m}(\gamma^{1/2}z)^{m}\right)\exp\left(\sum_{m>0}\frac{p^{m}}{1-p^{m}}\alpha_{m}(\gamma^{1/2}z)^{-m}\right).$
@@
$C,\gamma^{1/2}\ :\ \mbox{central},$ (2.2)
$[\alpha_{m},\alpha_{n}]=-\frac{\kappa_{m}}{m}(\gamma^{m}-\gamma^{-m})\gamma^{-m}\frac{1-p^{m}}{1-p^{*m}}\delta_{m+n,0},$ (2.3)
$[\alpha_{m},x^{+}(z)]=-\frac{\kappa_{m}}{m}\frac{1-p^{m}}{1-p^{*m}}\gamma^{-m}z^{m}x^{+}(z)\quad(m\not=0).$ (2.4)
$[\alpha_{m},x^{-}(z)]=\frac{\kappa_{m}}{m}z^{m}x^{-}(z)\quad(m\not=0),$ (2.5)
$[x^{+}(z),x^{-}(w)]=-\frac{(1-t_{1})(1-t_{2})}{(1-\hbar)}\left(\delta(\gamma^{-1}z/w)\psi^{+}(\gamma^{1/2}w)-\delta(\gamma z/w)\psi^{-}(\gamma^{-1/2}w)\right),$ (2.6)
$z^{3}G^{+}(w/z)g(w/z;p^{*})x^{+}(z)x^{+}(w)=-w^{3}G^{+}(z/w)g(z/w;p^{*})x^{+}(w)x^{+}(z),$ (2.7)
$z^{3}G^{-}(w/z)g(w/z;p)^{-1}x^{-}(z)x^{-}(w)=-w^{3}G^{-}(z/w)g(z/w;p)^{-1}x^{-}(w)x^{-}(z),$ (2.8)
$g(w/z;p^{*})g(u/w;p^{*})g(u/z;p^{*})\left(\frac{w}{u}+\frac{w}{z}-\frac{z}{w}-\frac{u}{w}\right)x^{+}(z)x^{+}(w)x^{+}(u){}$
+permutations in $+\mbox{permutations in $z,w,u$}=0,$ (2.9)
$g(w/z;p)^{-1}g(u/w;p)^{-1}g(u/z;p)^{-1}\left(\frac{w}{u}+\frac{w}{z}-\frac{z}{w}-\frac{u}{w}\right)x^{-}(z)x^{-}(w)x^{-}(u){}$
+permutations in $+\mbox{permutations in $z,w,u$}=0,$ (2.10)
@@
for $s=p,p^{*}$.
@@
Remark. For ${\cal U}$-modules, on which the central element $\gamma^{1/2}$ takes a complex value, we regard $p,p^{*}=p\gamma^{-2}$ as a generic complex number with $|p|<1,|p^{*}|<1$. Then we have
@@
Then one can rewrite, for example, (2.7)-(2.8) in the sense of analytic continuation as follows.
@@
2.2 Hopf algebroid structure

### v2

# Month $\rvsdateMonth Day, Year$

@@ Abstract.
We propose new vertex operators, both the type I and the type II dual, of the elliptic quantum toroidal algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ by combining representations of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ and the notions of the elliptic stable envelopes for the instanton moduli space ${\mathcal{M}}(n,r)$. The vertex operators reproduce the shuffle product formula of the elliptic stable envelopes by their composition. We also show that the vacuum expectation value of a composition of the vertex operators gives a correct formula of the $\mathrm{K}$-theoretic vertex function for ${\mathcal{M}}(n,r)$. We then derive exchange relations among the vertex operators and construct a $L$-operator satisfying the $RLL=LLR^{*}$ relation with $R$ and $R^{*}$ being elliptic dynamical instanton $R$-matrices defined as transition matrices of the elliptic stable envelopes. Assuming a universal form of $L$, we define a comultiplication $\Delta$ in terms of it. It turns out that the new vertex operators are intertwining operators of the $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$-modules w.r.t $\Delta$.
Key words and phrases:
Elliptic Quantum Group, Toroidal algebra, Quiver Variety, Elliptic Stable Envelope, K-theory, Vertex Operator, Vertex Function
2020 Mathematics Subject Classification:
Primary 17B37; Secondary 14N10, 81R10
1. Introduction
@@
The aim of this paper is to present a new formulation of the vertex operators of the elliptic quantum toroidal algebra (EQTA) $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ by combining its representations and the notions of the elliptic stable envelopes for the instanton moduli space ${\mathcal{M}}(n,r)$.
[…]
@@
The notion of stable envelopes was initiated in [35, 41]. They provide a good basis of the equivariant cohomology and the $\mathrm{K}$-theory of Nakajima quiver varieties $X$ [37, 38] and have nice applications to enumerative geometry, geometric representation theory and quantum integrable systems. In particular, the transition matrices between stable envelopes defined for different chambers give the $R$-matrices satisfying the Yang-Baxter equation (YBE). Regarding such $R$-matrix as universal one i.e. a $L$-operator which generate a quantum group, the transition property can be understood as a Gauss decomposition of the $L$-operator in terms of the stable envelopes. Hence the stable envelopes provide a new geometric formulation of quantum groups as well as quantum integrable systems associated with the quiver varieties. The elliptic version of stable envelopes (ESE) [1] for the equivariant elliptic cohomology $\mathrm{E}_{T}(X)$ is more fascinating. They depend on both the equivariant and the Kähler parameters and provide the elliptic dynamical $R$-matrix as their transition matrix. There these two parameters play a role of the spectral and the dynamical parameters, respectively.
@@
An typical example is the case $X=T^{*}fl$, the cotangent bundle to the partial flag variety. The ESE’s for $\mathrm{E}_{T}(T^{*}fl)$ have been identified with the elliptic weight functions of the $A$ type appearing in the elliptic hypergeometric integral solutions of the elliptic $q$-KZ equation[28, 29, 43] and their transition matrices have turned out to coincide with the known elliptic dynamical $R$-matrices of type AA[17, 12, 18, 24]. In [28, 29], such identification was carried out by using the vertex operators defined as intertwining operators of the $U_{q,p}(\widehat{\mathfrak{sl}}_{N})$-modules w.r.t the standard comultiplication given in terms of the $L$-operator instead of the Drinfeld comultiplication. Basically such vertex operators are operator valued integrals, the so called screened vertex operators, possessing basic elliptic weight functions i.e. basic elliptic stable envelopes, as their integration kernels. Then a composition of them produces more general ESEs corresponding to general configurations in $T^{*}fl$ via the shuffle product formula[29]. Such structure can indeed be regarded as an inheritance from the $L$-operator of $U_{q,p}(\widehat{\mathfrak{sl}}_{N})$, whose Gauss components are given by the half currents, i.e. integrals of the elliptic currents of $U_{q,p}(\widehat{\mathfrak{sl}}_{N})$ with the same basic elliptic weight functions as integration kernels[22]. This is consistent to a picture that the stable envelopes give a half of the Gauss decomposition of the $L$-operators.
@@
In addition, for the same case $X=T^{*}fl$, it has been shown that the vertex operators yield a $\mathrm{K}$}-theoretic vertex function[41, 42, 33] as a vacuum expectation value of a composition of the vertex operators[30]. The vertex function is a generating function of the counting of quasi maps from ${\mathbb{P}}^{1}$ to $X$. It is given by a certain $q$-hypergeometric integral containing the elliptic weight function inside the integral and satisfies certain difference equations. The monodromy matrix of it is then given by a transition matrix of the weight functions. In [1], such elliptic weight functions are called the pole subtraction elliptic stable envelopes.
[…]
@@
This paper is organized as follows. Sec.2 is a review of the elliptic quantum toroidal algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ including the Hopf algebroid structure and representations of the level $(1,N)$ and $(0,1)$. In Sec.3, we summarize known results on the ESE for $\mathrm{E}_{T}({\mathcal{M}}(n,r))$. In Sec.4, we propose a new construction of the type I and the type II dual vertex operators of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ and check the shuffle product formula of ESE’s. Sec.5 is devoted to a study of the $\mathrm{K}$-theoretic vertex functions for ${\mathcal{M}}(n,r)$. In Sec.6, we derive exchange relations of the vertex operators. Finally, in Sec.7, we construct a $L$-operator $L^{+}(u)$, which satisfies the $RLL=LLR^{*}$ relation. The exchange relations between the vertex operators and $L^{+}(u)$ are given. Defining a new comultiplication $\Delta$, we show that our vertex operators are intertwining operators of the $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$-modules w.r.t. $\Delta$. For comparison, we list in Appendix A a summary of the vertex operators w.r.t. the Drinfeld comultiplication $\Delta^{D}$ obtained in [32].
2. Elliptic Quantum Toroidal Algebra ${U}_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$
@@
We review the elliptic quantum toriodal algebra $\mathcal{U}_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ introduced in [32]. Parameters $t_{1},t_{2},\hbar=t_{1}t_{2}$ in this paper correspond to $q^{-1},t,t/q$ in [32], respectively. We assume $t_{1},t_{2},p$ are generic complex numbers satisfying $|t_{1}|,|t_{2}|,|p|<1$.
2.1. Definition
@@
Let us consider the Heisenberg algebras generated by $c,\ \Lambda_{0},\ c^{\perp}$, $\Lambda_{0}^{\perp},\ h,\ \alpha,\ P,\ Q$ satisfying the commutation relations
$[c,\Lambda_{0}]=1=[c^{\perp},\Lambda_{0}^{\perp}]=[h,\alpha]=[P,Q],$ (2.1)
@@
the others are zero. We set $\gamma=\hbar^{c^{\perp}/2},\ C=\hbar^{c/2},\ \mathfrak{z}^{*}=\hbar^{P}$ and $\mathfrak{z}=\mathfrak{z}^{*}C^{-2}=\hbar^{P-c}$. We call $\mathfrak{z}^{*}$ the dynamical parameter. Let $\mathbb{F}$} be the field of meromorphic functions of $\mathfrak{z}$ and $\mathfrak{z}^{*}$. We have
$g(\mathfrak{z},\mathfrak{z}^{*})e^{-\Lambda_{0}-Q}=e^{-\Lambda_{0}-Q}g(\mathfrak{z},\mathfrak{z}^{*}\hbar^{-1}),\ \ g(\mathfrak{z},\mathfrak{z}^{*})e^{\Lambda_{0}}=e^{\Lambda_{0}}g(\mathfrak{z}\hbar^{-1},\mathfrak{z}^{*})\$
for $\forall g(\mathfrak{z},\mathfrak{z}^{*})\in{\mathbb{F}}$. Set
@@
[…]
@@
$\psi^{+}(z):=C\exp\left(-\sum_{m>0}\frac{p^{m}}{1-p^{m}}\alpha_{-m}(\gamma^{-1/2}z)^{m}\right){}$
$\qquad\qquad\quad\times\exp\left(\sum_{m>0}\frac{1}{1-p^{m}}\alpha_{m}(\gamma^{-1/2}z)^{-m}\right),$
$\psi^{-}(z):=C^{-1}\exp\left(-\sum_{m>0}\frac{1}{1-p^{m}}\alpha_{-m}(\gamma^{1/2}z)^{m}\right){}$
$\qquad\qquad\quad\times\exp\left(\sum_{m>0}\frac{p^{m}}{1-p^{m}}\alpha_{m}(\gamma^{1/2}z)^{-m}\right).$
@@
$\hskip-28.45274pt\gamma^{1/2},\ C\ :\ \mbox{central},$
$\hskip-28.45274pt[\alpha_{m},\alpha_{n}]=-\frac{\kappa_{m}}{m}(\gamma^{m}-\gamma^{-m})\gamma^{-m}\frac{1-p^{m}}{1-p^{*m}}\delta_{m+n,0},$
$\hskip-28.45274pt[\alpha_{m},x^{+}(z)]=-\frac{\kappa_{m}}{m}\frac{1-p^{m}}{1-p^{*m}}\gamma^{-m}z^{m}x^{+}(z)\quad(m\not=0).$
$\hskip-28.45274pt[\alpha_{m},x^{-}(z)]=\frac{\kappa_{m}}{m}z^{m}x^{-}(z)\quad(m\not=0),$
$\hskip-28.45274pt[x^{+}(z),x^{-}(w)]=-\frac{(1-t_{1})(1-t_{2})}{(1-\hbar)}{}$
$\qquad\qquad\quad\times\left(\delta(\gamma^{-1}z/w)\psi^{+}(\gamma^{1/2}w)-\delta(\gamma z/w)\psi^{-}(\gamma^{-1/2}w)\right),$
$\hskip-28.45274ptz^{3}G^{+}(w/z)g(w/z;p^{*})x^{+}(z)x^{+}(w)=-w^{3}G^{+}(z/w)g(z/w;p^{*})x^{+}(w)x^{+}(z),$
$\hskip-28.45274ptz^{3}G^{-}(w/z)g(w/z;p)^{-1}x^{-}(z)x^{-}(w)=-w^{3}G^{-}(z/w)g(z/w;p)^{-1}x^{-}(w)x^{-}(z),$
$\hskip-28.45274ptg(w/z;p^{*})g(u/w;p^{*})g(u/z;p^{*})\left(\frac{w}{u}+\frac{w}{z}-\frac{z}{w}-\frac{u}{w}\right)x^{+}(z)x^{+}(w)x^{+}(u){}$
$+\mbox{permutations in $z,w,u$}=0,$
$\hskip-28.45274ptg(w/z;p)^{-1}g(u/w;p)^{-1}g(u/z;p)^{-1}\left(\frac{w}{u}+\frac{w}{z}-\frac{z}{w}-\frac{u}{w}\right)x^{-}(z)x^{-}(w)x^{-}(u){}$
$+\mbox{permutations in $z,w,u$}=0,$
@@
for $s=p,p^{*}$. The dynamical parameter $\mathfrak{z}^{*}$ commutes with $\alpha_{m}$ and $x^{\pm}_{n}$.
@@
Remark. For $\mathcal{U}$}-modules, on which the central element $\gamma^{1/2}$ takes a complex value, we regard $p,p^{*}=p\gamma^{-2}$ as generic complex numbers with $|p|<1,|p^{*}|<1$. Then we have
@@
Then one can rewrite, for example the relations for $x^{+}(z),x^{-}(z)$ in the sense of analytic continuation as follows.
@@
2.2. Hopf algebroid structure